# This work was supported by project from DOE-CESER

### v1

@@ Abstract
The power grid is experiencing a multi-fold transformation while the global climate evolves with record-breaking extreme temperatures during heat domes, polar vortexes, and severe ice. Over the decades, these extreme temperature events have increased in frequency, duration, and intensity. The power grid infrastructure is geographically spread over thousands of square miles with millions of small and large components, and the impact of extreme temperature operations on the grid infrastructure needs to be researched further. This paper reviews academic literature, standards, industry articles, and federal reports to identify the impacts of heat domes, polar vortexes, and icing on all the T&D grid equipment, including substations (assets owned and operated by the utilities and independent system operators). This paper classifies the equipment into primary and auxiliary equipment and determines its vulnerability to extreme temperatures for a deeper analysis of a more critical and vulnerable set of grid equipment. For each equipment under consideration, its fundamental role in the system, the impact of extreme temperatures on its operation, available monitoring, and mitigation of these impacts are discussed. The paper develops insights on standards readiness and identifies gaps concerning extreme temperature definitions. The paper also develops summary tables to identify the critical failure modes for each type of equipment, failure influence diagrams, and cascading influence diagrams to highlight and aid in translating the equipment vulnerability information into power grid contingency definitions that need to be considered in grid planning and operations.
@@ I Introduction
The electric power grid is one of the largest man-made infrastructures reliably operating and energizing the world [1]. The power grid is a largely meshed power transmission system and a generally radial power distribution system spread over thousands of square miles. The transmission and distribution (T&D) systems have equipment to carry power and auxiliary equipment and components for measurement, communication, and control. However, over the recent decades, there has been a significant shift in the operating conditions of the T&D grid, mainly due to climate change [2]. The T&D grid infrastructure consists of hundreds of small and large components and systems that work in a coordinated manner to ensure a safe and reliable system operation [3]. These components range from large outdoor high voltage, high power transformers to small outdoor components like vibration dampers to large and small indoor equipment [4]. These components are experiencing operating conditions that predict when they are installed and commissioned [2]. The power grid equipment is typically designed to last for decades as frequent maintenance and repairs are very expensive [5]. However, the expected operating conditions (including projections) may differ when these were designed/commissioned. Therefore, it is essential to understand the present conditions and vulnerability to extreme operating conditions.
@@
Figure 1: Inter-dependencies of critical infrastructure on electricity [6]
Figure 2: Map of high-voltage transmission lines in U.S. [7]
Most of the grid equipment is above ground and exposed to weather and the elements. Fig. 2 depicts the density of high-voltage transmission lines in the U.S., which does not include the distribution lines. As climate change continues to progress, the components of the electrical system in the U.S. are subjected to changing weather conditions. Two weather events of interest are heat domes and polar vortexes. These extremely hot and extreme cold events are projected to threaten power grid equipment more and more as average global temperatures rise [8, 9, 10, 11, 12]. The electric power grid is a central critical infrastructure with interdependency with other critical infrastructures and is essential for all other critical infrastructures to function, as seen from Figure 1. Overall, it is essential to capture the impact of climate change on the power grid and mainly look at some extreme climate change scenarios.
I-A CLIMATE CHANGE TRENDS OVER THE RECENT YEARS
@@ I-A1 EXTREME HEAT
Over the recent years, the overall average temperatures across the globe have been increasing, leading to an increased number of heat waves for a longer duration [10].
(a) Heat wave occurrences and intensity in the United States by decade, $1961$-$2019$.
(b) Heat wave season duration in the United States by decade, $1961$-$2019$.
Figure 3: Heat wave characteristics in the U.S. over the past few decades [13]
Unlike heat waves, the temperatures can increase 30 degrees above the nominal temperature [deadly heat dome]. This can have a significant impact on the operating conditions of the power grid, making the power grid extremely vulnerable during such conditions. The most recent extreme heat dome was in June 2021 in the Pacific Northwest Region of the U.S. [14].
@@
Heat domes are weather events in which a high-pressure atmosphere pushes warm air toward the Earth’s surface. These conditions prevent the formation of clouds in the area, allowing sunlight to heat the air near the ground further. Heat domes cause extremely high temperatures that can last for days [16] in regions that cover multiple states (Fig. 4). Historical and projected data suggest that high temperatures caused by heat dome-like events are increasing in intensity at a higher rate than the average global temperatures [10]. One projection anticipates that the probability of extreme high-temperature events will grow at an increasing, non-linear rate as the global average temperature rises (Fig. 5) [11].
@@ I-A2 EXTREME COLD
Extreme heat caused by heat domes or heat waves is not the only concern for the power grid. The overall increase in global temperatures is causing disrupted polar vortexes, spreading the reach of polar vortexes closer to the southern states in the U.S. [12]. The polar vortex is a region of winds that circulate from west to east in the stratosphere above the Arctic. When the vortex weakens, it can move south from the pole as warmer air moves north (Fig. 6) [17, 12]. This event can send cold Arctic air into the U.S. as far south as Florida and Texas [17, 18] and result in freezing temperatures that last for days [12]. Climate change, resulting in warmer temperatures across the Arctic, causes more polar vortex disruptions, bringing these events to the U.S. with greater frequency and intensity [12].
Disrupted polar vortexes cause extreme cold conditions, leading to unexpected sudden drops in temperatures that the power grid is exposed to. The Arctic polar vortex affects the U.S. states and power grid, and with the increase in the number and intensity of polar vortexes, it is important to evaluate their impacts on the power grid.
Figure 7: Illustration of a stable and unstable polar vertex [12].
Fig. 7 shows the critical differences between a stable and disrupted polar vortex. Since the polar vortexes impact the air temperature significantly, much of the power grid equipment gets exposed to extremely cold temperatures, leading to a concern. The disrupted polar vortexes can trigger snow storms in regions that are not expected to have extreme cold temperatures.
I-B ORGANIZATION OF THE PAPER
The paper is organized to enable a better articulation and readability of the equipment impacts due to extreme heat and extreme cold operating conditions. The paper is divided into various sections, Section II describes some of the whether related events in United States. Section III presents the different equipment from transmission, distribution, and substation; and their vulnerability to extreme temperatures which is then used to identify the vulnerable equipment to study in this work. Sections IV, V, VI discusses the impact of climate change, methods to measure and mitigate the impacts on different power grid equipment. It is to be noted that certain power grid equipment are placed in transmission, distribution, and substation locations. For such equipment, their relevant information is placed in a single section where the failure impact is prevalent. For example, the fuses subsection is in the distribution system; the transformer are in the substation section; insulators and bushings are in the transmission section, etc. Section VII provides the summary of standards for different equipment on the power grid considering extreme temperatures due to climate change. Section VIII present the summary of previously mentioned sections in a consolidated manner, this includes 1) direct and indirect failure modes in different equipment due to climate change, 2) summary of standards considered in this work, 3) operational temperature (from standards) violations for different equipment across various US geographical locations considering abnormal climate change scenarios, and 4) inter-dependency and cascading behaviors of power grid equipment consider failure modes due to extreme temperatures. Section IX presents the brief background, conclusion, and future work.
II EVENTS DUE TO EXTREME TEMPERATURES ON POWER GRIDS
@@ II-A Hot temperature-related events
Heat domes and heat dome-like weather have long been sources of power outages in the U.S. Below are some reviewed causes, details, and effects of some of the most prominent heat-related outages starting from 1996.
@@ II-A1 August 10, 1996, West Coast
II-A2 June 10-11, 1999, Ohio, Michigan, Indiana
High temperatures that lingered for multiple days in Ohio, Michigan, and Indiana resulted in heavy transmission loading, exacerbated by equipment outages around Northwest Ohio. American Electric Power (AEP) transferred 4,270 MW of power to neighboring utilities when low voltage conditions warranted transmission loading relief. 175 MW from AEP to First Energy and 308 MW from Virginia Power to ComEd had to be curtailed. A 345 kV circuit and a 345/138 kV transformer went out due to a trip from a fault and the misoperation of a breaker.
@@ II-B Cold temperature-related events
Equipment failures and power outages in winter months have frequently been caused by polar vortexes’ severe cold and icing conditions. Some of the significant disturbances and their causes are reviewed below.
II-B1 December 4-5, 2002, North Carolina
An ice storm in North Carolina coated the state in 0.75-1 inch of ice. This ice built up on power lines, power lines carrying equipment, and trees. Many trees still had their leaves because the storm occurred early in the winter. This extra surface area allowed more ice to accumulate on the trees, which caused a great number of them to fall onto power lines. North Carolina suffered around 362,000 outages, and its three principal utilities experienced many equipment failures. Across these utilities 4,523 poles, 5,587 cross-arms, and 40,760 insulators were damaged. Additionally, 4,503 transformers and 116,048 fuses were replaced [22].
II-B2 February 1-5, 2011, Southwest
@@
II-B3 October 29-30, 2011, Northeast
A snowstorm in the Northeast caused significant power outages that lasted up to a week. In total, seventy-four transmission lines had outages, six caused by direct snow and ice accumulation on the equipment. Transmission failures caused forty-four transmission substations to go offline as well, and 130,000 customers across 6 states lost power. Trees that fell due to snow and ice build-up also brought down or shortened many distribution lines. Many more distribution lines were damaged than transmission lines because less clearance was kept between lower voltage lines and trees.
Additionally, a 115 kV power line was taken offline for 5 hours because a relay was tripped incorrectly, although no customers were affected. Another 115 kV line was taken out of service when a circuit breaker failed to operate. The circuit breaker was stuck and couldn’t isolate a faulted line, and this caused 4,900 consumers to lose power. An additional 3,800 lost power due to unknown failures that could result from trees touching lines or icing across insulators [24].
II-B4 February 2021, Texas
In 2021 an unusually long and cold polar vortex hit Texas. The cold temperatures caused icing to form on wind turbine blades and in their gearboxes, and there were issues with natural gas supply to electrical generation plants. While energy demand increased as customers turned their heaters up, large generators started to go offline unexpectedly. Weather-related outages increased from 15,000 MW to 30,000 MW in 24 hours, and transmission and substation outages caused 1900 MW of solar and wind generation to go offline. The frequency began to drop at the end of February $14^{th}$ because the load was reaching the maximum capacity, so generators began shedding load on the $15^{th}$. At one point, the frequency dropped to a near-critical point but recovered before a complete blackout occurred [25]. Overall, 111 people lost their lives, 4.5 million homes lost power, and \$130 billion of losses and property damage occurred as a result of the event [26].
@@ III SCOPING STUDY
The power grid infrastructure is a unique combination of legacy devices, equipment, constant upgrades, and power grid expansion. This makes several parts of the power grid vulnerable to climate change. The power grid also comprises many kinds of equipment and systems. In this work, the power grid equipment has been categorized in such a way as to make the analysis meaningful and manageable. Tab. II presents the scoping of different transmission, distribution, and substation equipment. This is achieved by analyzing two criteria - the criticality of equipment during transients and the vulnerability of the equipment due to extreme temperature. For example, if equipment is vulnerable to extreme temperatures in this work, we analyze the impact of equipment failure on criticality. From Tab. II, it can be observed that insulators are an example that is scoped in, whereas vibration dampers are not scoped in.
@@
TABLE I: Scoping of transmission, distribution, and substation equipment based on criticality and vulnerability to extreme temperatures.
@@
S.No Equipment Funtionality Category Criticality during transients Vulnerability to extreme temperature Scope (in/out)
1 Vibration dampers Prevent mechanical vibrations Performance Low Minimal Out
@@
2 Insulators Insulate towers and conductors Protection High Minimal In
3 Arcing horns Flashover protection Protection High Null Out
@@
4 Optical Ground Wire (OPGW) Lightening protection and telecommunication Communication High Null Out
5 Conductor Power transmission Operation High High In
6 Corona ring Avoid Corona discharge Performance Moderate Moderate In
@@
7 Jumper wire Connector between conductors Performance Moderate Null Out
8 Spacers Spacing between conductors Protection Moderate Null Out
9 Transformers Measurement and conversion of kV levels Operation High High In
@@
10 Riser Connect conduit from ground to overheadlines via potheads Operation Moderate Null Out
11 Circuit breakers (Air, oil, SF6, vacuum) Protection Protection and control High High In
12 Batteries Energy storage Storage High High In
@@
13 Bus support insulators / Bushings Isolate bus bar switches and support structures Protection High Minimal In
14 Capacitor bank Control voltage level Regulation Moderate High In
15 Circuit switchers Protection Protection Moderate Moderate In
16 Voltage regulator Voltage regulation Regulation High High In
@@
17 Distribution busbars steel structure array of switches used to route power out of a substation Operation Moderate Moderate In
@@
18 Frequency changers changes power of an alternating current system from one frequency to one or more different frequencies Regulation High High In
19 Grounding resistors Limit transient currents to protect and operate Protection High High In
20 Fuses Protection Protection High High In
21 Lightening arresters Protection Protection High High In
22 Relays Protection Protection and control High High In
23 Shunt reactors Control voltage level Regulation High High In
24 Suspension insulators Protection Protection High Moderate In
25 Distribution Transformers Conversion Operation High High In
26 Underground Cables Power transmission Operation High High In
@@
27 OH-Line Insulators (distribution and medium voltages) Protection Protection High High In
28 Power Line Carrier Equipment Communication Communication High High In
29 CTs Measurement and control Measurement High High In

### v2

@@ Abstract
The power grid is experiencing a multi-fold transformation while the global climate evolves with record-breaking extreme temperatures during heat domes, polar vortices, and severe ice. Over the decades, these extreme temperature events have increased in frequency, duration, and intensity. The power grid infrastructure is geographically spread over thousands of square miles with millions of small and large components, and the impact of extreme temperature operations on the grid infrastructure needs to be researched further. This paper reviews academic literature, standards, industry articles, and federal reports to identify the impacts of heat domes, polar vortices, and icing on all the T&D grid equipment, including substations (assets owned and operated by the utilities and independent system operators). This paper classifies the equipment into primary and auxiliary equipment and determines its vulnerability to extreme temperatures for a deeper analysis of a more critical and vulnerable set of grid equipment. For each equipment under consideration, its fundamental role in the system, the impact of extreme temperatures on its operation, available monitoring, and mitigation of these impacts are discussed. The paper develops insights on standards readiness and identifies gaps concerning extreme temperature definitions. The paper also develops summary tables to identify the critical failure modes for each type of equipment, failure influence diagrams, and cascading influence diagrams to highlight and aid in translating the equipment vulnerability information into power grid contingency definitions that need to be considered in grid planning and operations.
@@ I Introduction
Electric power grids, consisting of transmission and distribution (T&D) systems spread over thousands of square miles, are among the largest infrastructures on Earth [1]. The transmission system is largely meshed, and the distribution system is generally radial. Each system has equipment to carry power, as well as auxiliary equipment and components for measurement, communication, and control. Together, the grid contains hundreds of small and large components and systems that work in a coordinated manner to ensure safe and reliable grid operation [2]. These components range from large outdoor high-voltage, high-power transformers, to small outdoor components like vibration dampers, to large and small indoor equipment [3]. Power grid equipment is typically designed to last for decades, because more frequent maintenance and repairs would be very expensive [4]. However, equipment lifespans are estimated on design basis assumptions for historically stable operating conditions. There has been a significant shift in the operating conditions of the T&D grid in recent decades—–in particular, higher incidence of extreme weather related to climate change [5]. Therefore, it is essential to understand the how operating conditions for T&D equipment may differ from when this equipment was designed or commissioned and assess its vulnerability to extreme operating conditions.
@@
Figure 1: Interdependence of critical infrastructure on electricity [6]
Figure 2: Map of high-voltage substations (light orange bubbles at terminals of the transmission lines in the map) and transmission lines in U.S. [7]
Importantly, the electric power grid is a central critical infrastructure; it is essential for all other critical infrastructures to function (Fig. 1). Fig. 2 shows the density of high-voltage substations and transmission lines in the U.S.; most T&D equipment is above ground and exposed to weather. Two weather events of interest are heat domes and polar vortices. These extreme hot and cold events are projected to threaten power grid equipment more and more as average global temperatures rise [8, 9, 10, 11, 12]. They are key areas of focus when considering future impacts to the grid, which in turn have compounded effects to other critical infrastructures.
I-A TEMPERATURE TRENDS AFFECTING POWER GRID EQUIPMENT
@@ I-A1 EXTREME HEAT
Average temperatures have increased in recent decades across the globe, leading to an increased number of heat waves, a longer average duration per heat wave, and a longer heat wave season (Fig. 3). Heat domes contribute strongly to the most extreme heat waves [10].
Figure 3: Heat wave characteristics in the U.S. over the past six decades [13]. Frequency is the average number of heat wave occurrences. Intensity is the average temperature (in °$\degree$°F) above local threshold during heat waves. Average duration is the average length of annual heat wave season in days.
Heat domes are weather events in which atmospheric currents trap warm air against the Earth’s surface. These conditions prevent cloud formation, allowing sunlight to further heat the air near the ground. Heat domes cause extremely high temperatures that can last for days [14] and cover multiple states (Fig. 4), significantly affecting operating conditions for T&D equipment and simultaneously increasing cooling-related electric demand. These factors make the power grid extremely vulnerable during such conditions.
The most recent extreme heat dome was in June 2021 in the United States’ Pacific Northwest region [15]. Historical and projected data suggest that high temperatures caused by heat dome-like events are increasing in intensity at a higher rate than the average global temperatures [10]. One projection anticipates that the probability of extreme high-temperature events will grow at an increasing, non-linear rate as the global average temperature rises (Fig. 5) [11].
@@ I-A2 EXTREME COLD
The overall increase in global temperatures is causing extreme cold events as well as extreme heat events. Warmer temperatures disrupt polar vortices and cause them to reach as far as the southern states in the U.S. [12]. The polar vortex is a region of winds that circulate from west to east in the stratosphere above the Arctic. When the vortex weakens, it can move south from the pole as warmer air moves north (Fig. 6) [17, 12]. This pattern can send cold Arctic air into the U.S. as far south as Florida and Texas [17, 18] and result in freezing temperatures that last for days [12]. Warmer temperatures across the Arctic cause more polar vortex disruptions, bringing these events to the U.S. with greater frequency and intensity [12]. Fig. 7 shows the critical differences between a stable and disrupted polar vortex.
The extreme cold conditions caused by disrupted polar vortices can be accompanied by snow and ice storms in regions where these weather events are not regularly anticipated, challenging electric grid reliability even as heating-related electric demand reaches potentially unforeseen levels. The exposure of T&D equipment to colder air temperatures is also a cause for concern.
Figure 7: Stable and unstable polar vortices [12].
The paper is focuses on the impact of the equipment from a electrical function and does not explicitly focus on structural integrity that can arise due to corrosion in heat domes (it is a combination of temperature and humidity). While structural integrity related failures or degradation can impact the electrical function, this paper is a review of findings in the open literature on the electrical functionality that gets impacted during extreme heat and extreme cold related events.
I-B KEY CONTRIBUTIONS
The overarching contribution of this review paper is that it collates and combines a vast literature into a concise yet comprehensive document that enables the reader to understand and reflect on the key impacts due to extreme temperature operation of the power grid. The key contributions of the paper are as follows:
  1. 1.
    The paper provides a detailed review of acute failure modes, monitoring techniques, and mitigation approaches for all major and minor power grid equipment under extreme temperature operation. The power grid equipment spans across the transmission, distribution, and substation assets.
  2. 2.
    This paper consolidates the failure modes in a comprehensive manner for various power grid assets due to extreme heat and extreme cold scenarios.
  3. 3.
    The paper also consolidates the information from equipment standards to reflect the readiness of power grid standards to handle the extreme temperatures in the recent years. Furthermore, this paper also identifies the potential operation of power grid assets in non-compliance with standard specifications due to worsening climatic conditions such as heat wave in cold region, ice storm in a hot region, etc.
  4. 4.
    The paper provides two kinds of failure maps: influence maps that show the inter-dependency between different equipment failures that are caused due to extreme temperatures; and cascading failure maps that provides the details of how failure of a single critical equipment can cascade into failure of other system equipment.
I-C ORGANIZATION OF THE PAPER
The paper is organized as follows. Section II provides illustrative examples of extreme hot and cold weather events in the U.S. that caused failures of T&D equipment. Section III describes the methods of this study for choosing which classes of T&D equipment to evaluate and presents the results of that assessment. Sections IV, V, and VI discuss transmission, distribution, and substation equipment, respectively. For each class of equipment within those categories, these sections describe: 1) Impact of Extreme Temperature Operations, 2) methods to measure impacts of extreme weather, and 3)  methods to mitigate these impacts. Note that in cases where equipment fits into more than one category, this paper places it in the section where the impact of failure is most prevalent. For example, fuses are in the distribution section; transformers are in the substation section; insulators and bushings are in the transmission section; etc. Section VII summarizes standards related to extreme temperatures for different power grid equipment. Section VIII presents a consolidated discussion of equipment failure modes during extreme temperatures, how current and future temperatures may violate standards for T&D equipment, and how the interdependency of grid equipment may lead to cascading failures during extreme temperatures. Section IX presents a brief background, conclusion, and future work.
II CASE STUDIES: TEMPERATURE-RELATED POWER OUTAGES
@@ II-A Hot temperature-related events
Heat waves have caused many power outages in the U.S. Some of the most prominent heat-related outages, starting from 1996, are overviewed below.
@@ II-A1 August 10, 1996, West Coast
II-A2 June 10–11, 1999, Ohio, Michigan, Indiana
High temperatures that lingered for multiple days in Ohio, Michigan, and Indiana resulted in heavy transmission loading, exacerbated by equipment outages around Northwest Ohio. American Electric Power (AEP) transferred 4,270 MW of power to neighboring utilities when low voltage conditions warranted transmission loading relief. 175 MW from AEP to First Energy and 308 MW from Virginia Power to ComEd had to be curtailed. A 345 kV circuit and a 345/138 kV transformer went out because of a trip from a fault and the misoperation of a breaker.
@@ II-B Cold temperature-related events
Equipment failures and power outages in winter months have frequently been caused by polar vortices’ severe cold and icing conditions. Some of the most significant disturbances and their causes are reviewed below.
II-B1 December 4–5, 2002, North Carolina
An ice storm in North Carolina coated the state in 0.75–1 inch of ice. This ice built up on power lines, power-line-carrying equipment, and trees. Many trees still had their leaves, because the storm occurred early in the winter. This extra surface area allowed more ice to accumulate on the trees, which caused a great number of them to fall onto power lines. North Carolina suffered around 362,000 outages, and its three principal utilities experienced many equipment failures. Across these utilities, 4,523 poles, 5,587 cross-arms, and 40,760 insulators were damaged. Additionally, 4,503 transformers and 116,048 fuses were replaced [22].
II-B2 February 1–5, 2011, Southwest
@@
II-B3 October 29–30, 2011, Northeast
A snowstorm in the Northeast caused significant power outages that lasted up to a week. In total, 74 transmission lines had outages, six of which were caused by direct snow and ice accumulation on the equipment. Transmission failures caused forty-four transmission substations to go offline as well, and 130,000 customers across six states lost power. Trees that fell because of snow and ice build-up also brought down or shortened many distribution lines. Many more distribution lines were damaged than transmission lines because less clearance was kept between lower-voltage lines and trees.
Additionally, a 115 kV power line was taken offline for five hours because a relay was tripped incorrectly, although no customers were affected. Another 115 kV line was taken out of service when a circuit breaker failed to operate. The circuit breaker was stuck and could not isolate a faulted line, causing 4,900 consumers to lose power. An additional 3,800 consumers lost power because of unknown failures that could result from trees touching lines or icing across insulators [24].
II-B4 February 11–20 2021, Texas
In 2021, an unusually long and cold polar vortex hit Texas (”the Great Texas Freeze”). The cold temperatures caused icing to form on wind turbine blades and in their gearboxes, and there were issues with natural gas supply to electrical generation plants. While energy demand increased as customers turned their heaters up, large generators started to go offline unexpectedly. Weather-related outages increased from 15,000  MW to 30,000 MW in 24 hours, and transmission and substation outages caused 1900 MW of solar and wind generation to go offline. The frequency began to drop at the end of February 14 because the load was reaching the maximum capacity, and on February 15, generators began shedding load. At one point, the frequency dropped to a near-critical point but recovered before a complete blackout occurred [25]. Overall, 111 people lost their lives, 4.5 million homes lost power, and \$130 billion of losses and property damage occurred as a result of the event [26].
While these events do not necessarily highlight only T&D equipment failures, they involves lot of impacts on the power grid in general, however, this paper is an attempt to review the impacts of extreme heat and extreme cold related events on the operation of the various T&D grid equipment.
@@ III SCOPING STUDY
The power grid comprises many kinds of equipment and systems, including legacy devices as well as new upgrades and expansions. To determine which equipment is to be evaluated, this study analyzes equipment based on two criteria: its importance in grid operations and its vulnerability to extreme temperatures. The equipment that are scoped for a deeper literature review are identified.
While the scoping study mainly looked at the electrical behavior, there are other known vulnerabilities that directly impact the structural integrity of various equipment that can also lead to significant damage to the T&D grid infrastructure. However a followup study and its findings are planned to be reported in a future publication. The scope would be too large to address in a single publication.
Tab. II overviews different classes of transmission, distribution, and substation equipment and notes which types are in scope for this paper. The column “In Scope (Yes/No)” specifies if an equipment is considered in this paper. The decision “Yes” is selected in two scenarios. These scenarios are as follows: scenario 1 - either of the columns “Importance During Grid Events” or “Vulnerability to Extreme Temperature” is a “high”, scenario 2 - when both of the columns “Importance During Grid Events” and “Vulnerability to Extreme Temperature” are moderate. In any other combination of scenarios, the scope selected is a “No”. For example, insulators (high importance) are in scope, but vibration dampers (low importance, minimal temperature vulnerability) are not.
@@
TABLE I: Scoping of transmission, distribution, and substation equipment based on importance and vulnerability to extreme temperatures.
@@
S.No Equipment Functionality Category Importance During Grid Events Vulnerability to Extreme Temperature In Scope (Yes/No)
1 vibration dampers prevent mechanical vibrations performance low minimal no
@@
2 insulators insulate conductors from support structures protection high minimal yes
3 arcing horns flashover protection protection moderate null no
@@
4 optical ground wire lightning protection and telecommunication communication moderate null no
5 conductor power transmission operation high high yes
6 corona ring avoid corona discharge performance moderate moderate yes
@@
7 jumper wire connector between conductors performance high null yes
8 spacers spacing between conductors protection moderate null no
9 transformers measurement and conversion of voltage levels operation high high yes
@@
10 riser connect conduit from ground to overhead lines via potheads operation high null yes
11 circuit breakers (air, oil, SF6, vacuum) protection protection and control high high yes
12 batteries energy storage storage high high yes
@@
13 bus support insulators/bushings isolate bus bar switches and support structures protection high minimal yes
14 capacitor bank control voltage level regulation high high yes
15 circuit switchers protection protection moderate moderate yes
16 voltage regulator voltage regulation regulation high high yes
@@
17 distribution busbars steel structure array of switches used to route power out of a substation operation moderate moderate yes
@@
18 frequency changers changes power of an alternating current system from one frequency to one or more different frequencies regulation high high yes
19 grounding resistors limits transient currents protection high high yes
20 fuses protection protection high high yes
21 lightning arresters protection protection high high yes
22 relays Protection protection and control high high yes
23 shunt reactors control voltage level regulation high high yes
24 suspension insulators protection protection high moderate yes
25 distribution transformers conversion operation high high yes
26 underground cables power transmission operation high high yes
@@
27 Overhead line insulators (distribution and medium voltages) protection protection high high yes
28 power line carrier equipment communication communication high high yes
29 CTs measurement and control measurement high high yes